OpenStudy (anonymous):

how to find x for ln(3-x)=1 ??

OpenStudy (anonymous):

is it ln(3)-ln(x)=ln(1) 0-ln(x)=0 ln(x)=0

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

so how to do it @rizwan_uet

OpenStudy (anonymous):

should i take the base e to remove "ln "

OpenStudy (anonymous):

3-x=e^1 x=e^1-3 x=-0.281718 @rizwan_uet

OpenStudy (anonymous):

ln(3-x) = 1 = ln e ln3 - lnx = ln e ln(3/x) = ln e 3/x = e now you can find x

OpenStudy (anonymous):

x=3/e x=1.1036 if i substitute x inside ln(3-1.1036) not = to 1 @rizwan_uet

OpenStudy (anonymous):

x=0.6399

ganeshie8 (ganeshie8):

\( \ln(3-x)=1 \) \( 3-x=e^1 \) \( x=3-e \)

OpenStudy (anonymous):

Thanks @ganeshie8

ganeshie8 (ganeshie8):

np :) e is just a number, plug the value of e if u want.

OpenStudy (anonymous):

yeah now i know that e=2.7182 @ganeshie8

ganeshie8 (ganeshie8):

good :)